# Color it, Code it, Cancel it: $k$-local dynamical decoupling from classical additive codes

### v1

@@ VI Conclusion and outlook
Several interesting directions remain open for future investigation. Our current constructions are based on DD groups that are subgroups of the Pauli group, which naturally support a mapping to additive codes over $\mathbb{F}_{4}$. An important next step is to extend this framework to DD groups that are subgroups of the Clifford group [78, 79]. Such sequences could achieve symmetrization of the system dynamics [41, 78], and could become an additional tool for Hamiltonian simulation [29]. In fact, Ref. [79] is closely aligned in spirit with our work, leveraging the symmetries of specific spin–spin interactions to design efficient decoupling sequences based on single-qubit Clifford groups. Another promising direction is to explore the robustness of our DD sequences. While our methods assume idealized pulses, known DD sequences exist that are robust against pulse imperfections [80, 81, 11, 12]. Developing a systematic procedure to generalize our DD framework to make it more robust is a stimulating question. Finally, it remains an open challenge to integrate the hypergraph-coloring approach with the Uhrig sequence family [15, 17, 16], which is known to provide arbitrary-order noise suppression. However, the sequence length might increase significantly with increasing system size.
VII Acknowledgment
@@
We thank members of the Bosco, Rimbach-Russ, Terhal, and Vandersypen groups for valuable discussions. We are grateful to Pablo Cova Fariña for pointing out that the Kitaev honeycomb model can be simulated using a dynamical decoupling sequence of length $L=4$. This research was supported by the EU through the H2024 QLSI2 project, by the Army Research Office under Award Number: W911NF-23-1-0110, and by NCCR Spin (grant number 225153). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Bounded control sequence
In this appendix, we review the bounded-control scheme for dynamical decoupling sequences [31]. The key idea is that the control operations are generated by a set of generators of the decoupling group $\mathcal{G}$. The decoupling sequence is implemented by continuously and smoothly evolving between elements of $\mathcal{G}$ via these generators, rather than through instantaneous (bang-bang) pulses.
@@
We define $\mathrm{Cay}(\mathcal{G},\Gamma_{\mathcal{G}})$ as the Cayley graph of the group $\mathcal{G}$ with respect to the generating set $\Gamma_{\mathcal{G}}$ [31, 30]. The Cayley graph $\mathrm{Cay}(\mathcal{G},\Gamma_{\mathcal{G}})$ is a directed graph where each vertex corresponds to a group element $g_{i}\in\mathcal{G}$. There is a directed edge from vertex $g_{i}$ to vertex $g_{i+1}$, labeled by the generator $\gamma_{k}$ iff $g_{i+1}=\gamma_{k}g_{i}$.
@@ Appendix B Superconducting qubit (bounded control)
As a first step, we describe how to implement the control pulses associated with the generators of the decoupling group in the quotient graph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$. Each generator $\gamma_{i}\in\mathcal{G}$ is a Pauli string of the form
@@
where each element $\gamma_{p_{i}}]_{k}$ acts on the $k$-th color class. To lift this operator from the quotient graph back to the original graph, we impose that $\gamma_{p_{i}}]_{k}$ acts uniformly on all qubits belonging to the color class $c_{k}$. That is, for each qubit $q_{j}\in c_{k}$, the same single-qubit Pauli operator $\gamma_{p_{i}}]_{k}$ is applied.
@@
The decomposition ensures that each control pulse can be executed using local, parallel single-qubit operations in a way that respects the color partition of the interaction graph.
@@
which applies a global $Z$ operator to each color class. This operator commutes with all $Z$-type Pauli strings, meaning the corresponding control Hamiltonian does not introduce any unwanted interactions that would require further cancellation. Crucially, it anticommutes with any single-qubit $X$ or $Y$ operator acting on a colored qubit, thereby enabling universal decoupling of single-qubit errors. As a result, the total length of the bounded-control decoupling sequence becomes

### v2

@@ VII Conclusion and outlook
Several interesting directions remain open for future investigation. Our current constructions are based on DD groups that are subgroups of the Pauli group, which naturally support a mapping to additive codes over $\mathbb{F}_{4}$. An important next step is to extend this framework to DD groups that are subgroups of the Clifford group [read2025platonic, read2025factor]. Such sequences could achieve symmetrization of the system dynamics [zanardi1999symmetrizing, read2025platonic], and could become an additional tool for Hamiltonian simulation [Choi2020]. In fact, Ref. [read2025factor] is closely aligned in spirit with our work, leveraging the symmetries of specific spin–spin interactions to design efficient decoupling sequences based on single-qubit Clifford groups. Another natural and important question is to incorporate selective DD sequences into quantum gate operations. As demonstrated in our recent work [nguyen2025single], inserting an echo pulse within a quantum gate that selectively flips interaction terms can reduce gate infidelity by at least an order of magnitude. A systematic extension of our framework to general quantum operations may be possible by combining it with dynamically corrected gate protocols [Khodjasteh2009]. Finally, our present analysis is restricted to DD sequences with an explicit group structure and equally spaced pulses. However, more general DD protocols that lack a strict group structure and employ non-equidistant pulse timings have been shown to provide higher-order error suppression with fewer pulses [Uhrig2007, yang2008universality, wang2011protection, zhang2025learningsteerquantummanybody], as well as the ability to engineer effective interaction strengths beyond simple on–off control [hayes2014programmable, o2019hamiltonian, rajabi2019dynamical, Choi2020]. Extending our hypergraph-coloring–based approach to encompass these more general DD schemes remain an open question.
VIII Acknowledgment
We thank members of the Bosco, Rimbach-Russ, Terhal, and Vandersypen groups for valuable discussions. We are grateful to Pablo Cova Fariña for pointing out that the Kitaev honeycomb model can be simulated using a dynamical decoupling sequence of length $L=4$. We thank Yuning Zhang for helpful discussions on constructing robust dynamical decoupling sequences and for sharing his Mathematica code.
@@ VIII Acknowledgment
This research was supported by the EU through the H2024 QLSI2 project, by the Army Research Office under Award Number: W911NF-23-1-0110, and by NCCR Spin (grant number 225153). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Bounded control sequence
In this appendix, we review the bounded-control scheme for dynamical decoupling sequences [Viola2003]. The key idea is that the control operations are generated by a set of generators of the decoupling group $\mathcal{G}$. The decoupling sequence is implemented by continuously and smoothly evolving between elements of $\mathcal{G}$ via these generators, rather than through instantaneous (bang-bang) pulses.
@@ Appendix B Superconducting qubit (bounded control)
As a first step, we describe how to implement the control pulses associated with the generators of the decoupling group in the quotient hypergraph $I_{\rm Dev}^{{}^{\prime}}/C[I_{\rm Dev}]$. Each generator $\gamma_{i}\in\mathcal{G}$ is a Pauli string of the form
@@
where each element $\gamma_{p_{i}}]_{k}$ acts on the $k$-th color class. To lift this operator from the quotient hypergraph back to the original interaction hypergraph, we impose that $\gamma_{p_{i}}]_{k}$ acts uniformly on all qubits belonging to the color class $c_{k}$. That is, for each qubit $q_{j}\in c_{k}$, the same single-qubit Pauli operator $\gamma_{p_{i}}]_{k}$ is applied.
@@
The decomposition ensures that each control pulse can be executed using local, parallel single-qubit operations in a way that respects the color partition of the interaction hypergraph.
@@
which applies a global $Z$ operator to each color class. This operator commutes with all $Z$-type Pauli strings, meaning the corresponding control Hamiltonian does not introduce any unwanted interactions that would require further cancellation. Crucially, it anticommutes with any single-qubit $X$ or $Y$ operator acting on a color class, thereby enabling universal decoupling of single-qubit errors. As a result, the total length of the bounded-control decoupling sequence becomes
@@ Appendix D Simulating Kitaev’s honeycomb model with bounded control
Appendix E Sequence robustness under misaligned rotation axis
In Sec. VI of the main text, we assumed that the dominant pulse error arises from systematic over- or under-rotation and presented strategies for making DD sequences robust against such errors. In this appendix, we analyze the case in which the dominant pulse error instead originates from misalignment of the rotation axis, and we explain why the proposed strategies are ineffective against this type of pulse error.
As a concrete example, we consider a noisy implementation of the $(\tilde{X})_{\theta}$ pulse given by
$(\tilde{X})_{\theta}\equiv\exp\Big[-\frac{i}{2}\theta(\sqrt{1-\eta_{y}^{2}-\eta_{z}^{2}}X+\eta_{y}Y+\eta_{z}Z)\Big],$ (116)
where $\eta_{y},\eta_{z}\ll 1$ quantify small deviations of the rotation axis away from the ideal $X$ direction. As in the main text, we decompose the noisy pulse into an error operator $E[(X)_{\theta}]$ followed by an ideal $(X)_{\theta}$ pulse
$(\tilde{X})_{\theta}=(X)_{\theta}~E[(X)_{\theta}].$ (117)
To determine the explicit form of the error operator $E[(X)_{\theta}]$, we employ the following math trick. Consider
$\exp(i\lambda H)\exp[-i\lambda(H+\eta)]=F(\lambda),$ (118)
with $F(0)=I$. Differentiating both sides with respect to $\lambda$ yields
$\frac{dF(\lambda)}{d\lambda}$ $=\exp(i\lambda H)(-i\eta)\exp[-i\lambda(H+\eta)]$ (119a)
$=\exp(i\lambda H)(-i\eta)\exp(-i\lambda H),F(\lambda),$ (119b)
where, in going from the first to the second line, we have inserted the identity $I=e^{-i\lambda H}e^{i\lambda H}$ and used Eq. (118). To first order in $\eta||$, the solution at $\lambda=1$ is given by the first-order Magnus expansion,
$F(1)\approx\exp\Big[-i\int_{0}^{1}d\lambda\big(\exp[i\lambda H]\eta\exp[-i\lambda H]\big)\Big].$ (120)
Applying this result to the present case, with the identifications
$\exp[-i\lambda(H+\eta)]=(\tilde{X})_{\theta},\quad\exp[-i\lambda H]=(X)_{\theta},$ (121)
and $\theta=\pi$, we obtain the error operator
$E[(X)_{\pi}]\approx\exp\Big[-i\Big(-\eta_{y}Z+\eta_{z}Y\Big)\Big].$ (122)
Importantly, we find that the error operator associated with a noisy $(\tilde{X})_{-\pi}$ pulse is identical to that of a $(\tilde{X})_{\pi}$ pulse. To first order in the misalignment parameters $\eta$, we have
$E[(X)_{-\pi}]\approx E[(X)_{\pi}].$ (123)
This result can be directly verified using the identity
$I$ $=(\tilde{X})_{-\pi}(\tilde{X})_{\pi}$ (124a)
$\approx\exp\Big[i\big(-\eta_{y}Z+\eta_{z}Y\big)\Big]\exp\Big[-i\big(-\eta_{y}Z+\eta_{z}Y\big)\Big]$ (124b)
$=I,$ (124c)
where we have kept terms only to first order in $\eta$. The fact that $E[(X)_{-\pi}]$ coincides with $E[(X)_{\pi}]$ for axis-misalignment errors has important consequences. By contrast, for over- and under-rotation errors one finds $E[(X)_{-\pi}]=E[(X)_{\pi}]^{\dagger}$, a property that allows cumulative pulse errors $\tilde{E}$ (given in Eq. (82)) to be canceled by alternating between $(\pi)$ and $(-\pi)$ pulses, as discussed in Sec. VI.2.1. This cancellation mechanism, however, is absent for misaligned rotation axis errors. Consequently, alternating between $(\pi)$ and $(-\pi)$ pulses is ineffective in this case, and suppressing the accumulated pulse error instead requires choosing a Hamiltonian cycle such that the exponent on the right-hand side of Eq. (82) vanishes.
Another important distinction arises in the commutation properties of the pulse error $E[(X)_{\pi}]$ through the ideal pulse $(X)_{\pi}$. For misaligned rotation axis errors, we find
$(X)_{\pi}~E[(X)_{\pi}]=E[(X)_{\pi}]^{\dagger}(X)_{\pi}$ (125)
whereas for over- and under-rotation errors the corresponding relation is
$(X)_{\pi}~E[(X)_{\pi}]=E[(X)_{\pi}](X)_{\pi}.$ (126)
Owing to the property in Eq. (125), the mirror-sequence strategy introduced in Sec. VI.2.2 can in fact be used to suppress the cumulative pulse errors $\tilde{E}$ arising from axis misalignment. However, this modified commutation relation also leads to a crucial limitation. Repeating the reasoning following Eq. (93), one finds that the mirror sequence no longer cancels the residual Hamiltonian error $H_{\rm error}$. More explicitly, for misaligned rotation-axis errors, if the noisy DD sequence takes the form
$U_{\Delta}\tilde{p}_{L}\dots U_{\Delta}\tilde{p}_{1}=\Big(\prod_{i=1}^{L}\tilde{E}[p_{i}]\Big)e^{-i\Delta L(H_{\rm targ}+H_{\rm error})},$ (127)
then the corresponding mirror sequence necessarily has the structure
$\tilde{p}_{1}U_{\Delta}\dots\tilde{p}_{L}U_{\Delta}=e^{-i\Delta L(H_{\rm targ}-H_{\rm error})}\Big(\prod_{i=1}^{L}\tilde{E}[p_{i}]\Big)^{\dagger}.$ (128)
As a result, while the cumulative propagated pulse error $\Big(\prod_{i=1}^{L}\tilde{E}[p_{i}]\Big)$ cancel out, the residual Hamiltonian errors add up, preventing full error suppression.
In summary, for misaligned rotation-axis errors, the pulse-scheduling strategies presented in Sec. VI.2.1 are no longer effective. The mirror-sequence extension discussed in Sec. VI.2.2 can partially suppress pulse errors, but it does not eliminate them entirely. Consequently, an additional strategy is required to fully mitigate errors due to rotation-axis misalignment. We do not identify a simple solution within the scope of this work and therefore leave this problem for future investigation.